### v1

# Inflationary power spectrum from the Lanczos algorithm

Abstract
The generalized Lanczos algorithm can provide a universal method for constructing the wave function under the group structure of Hamiltonian. Based on this fact, we obtain an open two-mode squeezed state as the quantum origin for the curvature perturbation. In light of this wave function in the open system, we successfully develop a new method to calculate its corresponding power spectrum by using the Bogoliubov transformation. Unlike traditional approaches, we explicitly retain the Bogoliubov coefficients in terms of the squeezing amplitude $r_{k}$ and the squeezing rotation angle $\phi_{k}$. As a result, the power spectrum of the open two-mode squeezed state will match that of the Bunch-Davies vacuum numerically. Furthermore, the derivation of the open two-mode squeezed state relies on the second kind Meixner polynomial (equivalent to the generalized Lanczos algorithm) and the symmetry of the Hamiltonian. Therefore, our research may offer a new insight into the calculation of the correlation functions through a group-theoretic perspective.
I Introduction
[…]
The quantum origin of curvature perturbation has attracted much attention [13, 14]. It motivates us to use alternative methods from quantum information to evaluate the curvature perturbation. One interesting direction is to implement the concept of complexity to study the curvature perturbations. There are various kinds of complexities in quantum information, such as computational complexity [15], Krylov complexity [46], and etc. The computational complexity in inflation was examined in [17, 18, 19, 20, 21, 22, 23]. Among of these, Ref. [24] has studied the Krylov complexity in cosmology with non-trivial sound speed. However, those literatures studied the complexity for the closed system. But our universe is an open system, therefore, one must use the open system approach to study the Krylov complexity in inflation [25, 26]. In [25, 26], the authors have successfully adopted the two-mode squeezed state to study the Krylov complexity in inflation. As Ref. [47] mentioned, the amplification from the quantum level to the classical level in inflation is a quantum squeezing process. Thus, it is quite natural to use the two-mode squeezed state to evaluate the quantum fluctuations in inflation. The two-mode squeezed state has been broadly applied to the pre-inflationary phase [27, 28, 29, 30, 31, 32].
In this work, we will utilize the Bogoliubov transformation to establish a relationship between the two-mode squeezed state and the Bunch-Davies vacuum (BD). The Bogoliubov coefficients, denoted as $\alpha_{k}$ and $\beta_{k}$, are determined according to [33]. However, the wave function we employ is not the standard pure two-mode squeezed state; instead, we refer to it as the open two-mode squeezed state (OTMSS), which is derived from the second kind of Meixner polynomial (equivalent to the generalized Lanczos algorithm) and the symmetries of the Hamiltonian [26]. Following this, we then calculate the power spectrum of the OTMSS within the framework of the generalized Lanczos coefficient. And we find that the numerical results match those in the BD vacuum.
This work is organized as follows. Sec. II will give some foundations of curvature perturbation under the framework of single-field inflation. Sec. III describes the generalized Lanczos algorithm in inflation. In Sec. IV, we will derive the so-called OTMSS. Sec. V will show the explicit formula of $\alpha_{k}$ and $\beta_{k}$. Then, Sec. VI will obtain the numerics of $\phi_{k}$ and $r_{k}$. Based on this, Sec. VII will derive the power spectrum of OTMSS. Finally, the conclusion and discussion are drawn in Sec. VIII.
II Some foundation of inflationary perturbation
In this work, we will follow Ref. [34] to give some foundation for the cosmological perturbation. First, it is the spatially conformal flat Friedmann-Lemaitre-Robertson-Walker (FLRW) metric,
$ds^{2}=a^{2}(\eta)[-(1+\psi(\eta,\vec{x}))d\eta^{2}+(1-\psi(\eta,\vec{x}))d\vec{x}^{2}],$ (1)
where $\psi(\eta,\vec{x})$ is the perturbation of metric and $a(\eta)$ is the scale factor. Meanwhile, we can also define the perturbation of the inflaton as $\phi(x_{\mu})=\phi_{0}(\eta)+\delta\phi(x_{\mu})$, where $\phi_{0}(\eta)$ is the background field of inflaton and the $\delta\phi(x_{\mu})$ can be dubbed as the perturbation of inflaton. Being armed with these perturbations, we can obtain the generically quadratic action as follows,
[…]
where the Hubble rate $H=\dot{a}/a$ (denoting the derivation with the physical time), the conserved curvature perturbations is $\mathcal{R}=\psi+\frac{H}{\phi_{0}}\delta\phi$ at horizon crossing $(k=aH)$ and the Hubble flow expansion $\epsilon=-\dot{H}/H^{2}$, $z=\sqrt{2\epsilon}a$. Furthermore, we rewrite action (2) in terms of the Mukhanov-Sasaki variable $v=z\mathcal{R}$,
[…]
Applying the Euler-Lagrange equation, we obtain the equation of motion in the momentum space,
$v_{k}^{\prime\prime}+(k^{2}-\frac{z^{\prime\prime}}{z})v_{k}=0.$ (4)
One possible solution of Eq. (4) is the well-known BD vacuum,
$v_{BD}(\eta,\vec{k})=\frac{e^{-ik\eta}}{\sqrt{2k}}(1-\frac{i}{k\eta}).$ (5)
However, defining a privileged quantum vacuum state in curved spacetime is fundamentally problematic, as established by [35]. In this context, the BD vacuum emerges as a particular state that asymptotically approaches the Minkowski vacuum at short distances [36]. The complete vacuum characterization requires consideration of both the Poincaré patch of de Sitter spacetime (PdS) and Hadamard regularity conditions. Following standard cosmological practice, we maintain the BD vacuum assumption under PdS invariance and Hadamard conditions, consistent with foundational work [37, 38, 33].
From an operational perspective, establishing the BD vacuum requires two key elements: $(a)$. The initial conditions for Equation (4) in the sub-horizon regime preceding inflation. $(b)$. The harmonic approximation of (4) as a quantum harmonic oscillator.
The initial condition of Eq.(4) is that it can be approximated as a quantum harmonic oscillator in sub-horizon regime before inflation. These conditions demonstrate that the BD vacuum represents a highly specialized choice for inflationary quantum perturbations. We propose that a more fundamental vacuum state might exist, derivable through Hamiltonian symmetry principles and the Lanczos algorithm. This approach could reveal deeper structure in the quantum initial conditions of inflation.
III Lanczos algorithm in inflation
Recent studies [25, 39, 40, 26, 41] have demonstrated that the early universe behaved as a dissipative system, owing to both the violation of energy conservation [42] and the particle production during preheating [43, 44]. Motivated by these findings, we employ the Lanczos algorithm for open quantum systems to evaluate the quadratic action (3). This approach aligns with the generalized Lanczos algorithm framework developed in [45]. Following our previous conventions [25], we express the generalized Lanczos algorithm as
$\mathcal{L}|\mathcal{O}_{n})=-ic_{n}|\mathcal{O}_{n})+b_{n+1}|\mathcal{O}_{n+1})+b_{n}|\mathcal{O}_{n-1}),$ (6)
where the coefficients $c_{n}$ encode the open quantum system dynamics for a given Hamiltonian, while $b_{n}$ (the Lanczos coefficients) characterize the chaotic properties of the dynamical system. The orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$ spans the Krylov space, where it denotes the space of the two-mode squeezed state, and $\mathcal{L}$ denotes the Liouvillian superoperator – which can be represented in terms of creation and annihilation operators analogous to a Hamiltonian operator.
Following our previous work [25, 39, 40], we identify that:
  • The $c_{n}$ terms in (6) correspond to open system effects
  • The $b_{n}$ terms represent closed system dynamics
This formulation allows us to interpret $\mathcal{L}$ as an effective Hamiltonian that acts on two-mode quantum states, enabling explicit separation between open and closed system contributions. The action of $\mathcal{L}$ on suitable quantum states thus provides a mechanism to:
  1. 1.
    Isolate environmental decoherence effects (open system)
  2. 2.
    Preserve unitary evolution (closed system)
For comprehensive discussions of the Lanczos algorithm implementation and Krylov complexity in this context, we refer readers to [46].
[…]
$\pi(\eta,\vec{x})=\frac{\delta L}{\delta v^{\prime}(\eta,\vec{x})}=v^{\prime}-\frac{a^{\prime}}{a}v.$ (7)
Therefore, we could obtain the Hamiltonian
[…]
Utilizing the Fourier decomposition for $v$ and $\pi$,
$\hat{v}(\eta,\vec{x})=\int\frac{d^{3}k}{(2\pi)^{3/2}}\sqrt{\frac{1}{2k}}(\hat{c}_{-\vec{k}}^{\dagger}v_{\vec{k}}^{\ast}e^{-i\vec{k\cdot}\vec{x}}+{\hat{c}_{\vec{k}}}v_{\vec{k}}e^{i\vec{k\cdot}\vec{x}}),$ (9)
$\hat{\pi}(\eta,\vec{x})=i\int\frac{d^{3}k}{(2\pi)^{3/2}}\sqrt{\frac{k}{2}}(\hat{c}_{-\vec{k}}^{\dagger}u_{\vec{k}}^{\ast}e^{-i\vec{k\cdot}\vec{x}}-\hat{c}_{\vec{k}}u_{\vec{k}}e^{i\vec{k\cdot}\vec{x}}).$ (10)
where $\hat{c}_{-\vec{k}}^{\dagger}$ and $\hat{c}_{\vec{k}}$ denote the creation and annihilation operators, respectively, $u_{k}$ and $v_{k}$ are the function of conformal time $\eta$. Here, we keep two modes $(k,-k)$ since it is invariant under changing from $k$ to $-k$ in action (3) like in [36]. Therefore, the Hamiltonian (8) will become as follows
$\begin{split}\hat{H}=\int{d^{3}k}\hat{H}_{k}=&\int{d^{3}k}\bigg{[}k\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}+k\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}\\ &+i\frac{z{}^{\prime}}{z}\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}^{\dagger}-i\frac{z{}^{\prime}}{z}\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}\bigg{]}.\end{split}$ (11)
where have chosen appropriate normalizations for $v_{\vec{k}}$ and $u_{\vec{k}}$ and this Hamiltonian is of the Weyl-group structure [57]. Then, we can use this Hamiltonian (11) to act on the two-mode orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$, one can obtain
$\begin{split}\hat{H}|\mathcal{O}_{n})=(2n+1)k|\mathcal{O}_{n})+i\frac{z{}^{\prime}}{z}(n+1)|\mathcal{O}_{n+1})\\ -in\big{(}\frac{z{}^{\prime}}{z}\big{)}|\mathcal{O}_{n-1}).\end{split}$ (12)
where $b_{n}^{2}=n^{2}(\frac{z^{\prime}}{z})^{2}$, $c_{n}=i(2n+1)k$ and the prime symbol indicates differentiation with respect to conformal time. The formula of Lanczos coefficient clearly indicates that the early universe is a maximally chaotic, infinity dynamical system as discovered in [25, 39, 40, 26]. For convenience, we will define
$M_{P}^{2}|1-\mu_{1}^{2}|=\bigg{(}\frac{z^{\prime}}{z}\bigg{)}^{2},~{}~{}M_{P}\mu_{2}=k,$ (13)
where we have followed the notation of [25] and restored the unit of Planck mass, $\mu_{2}$ plays a role of dissipative coefficient and thus $b_{n}^{2}=|1-\mu_{1}^{2}|$ and $c_{n}=i(2n+1)\mu_{2}$. As previously mentioned, the $c_{n}$ part corresponds to the open system part of Hamiltonian (11), and the $b_{n}$ matches the closed system part of Hamiltonian (11). Thus, we could split the Hamiltonian into the open part and closed part as follows,
$\begin{split}&H_{open}=M_{P}\mu_{2}(\hat{c}_{-k}^{\dagger}\hat{c}_{-k}+\hat{c}_{k}\hat{c}_{k}^{\dagger}),\\ &H_{closed}=iM_{P}\sqrt{|1-\mu_{1}^{2}|}(\hat{c}_{k}^{\dagger}\hat{c}_{-k}^{\dagger}-\hat{c}_{k}\hat{c}_{-k}).\end{split}$ (14)
For any single field inflationary models, its Hamiltonian can be divided into two parts as shown in Eq. (14), where the information of various models are encoded in $\mu_{2}$ and $\mu_{1}$.
IV wave function
As Ref. [47] pointed out, the amplification from the quantum level to the classical level in inflation is a quantum squeezing process. Therefore, it is natural to use the two-mode squeezed state to evaluate the quantum fluctuations in inflation. However, in this work, we will implement a more generic function to replace this wave function. First and foremost, the so-called OTMSS via Ref. [25] is
$|\mathcal{O}(\eta))=e^{i\mathcal{L}\eta}|0,0\rangle_{-\vec{k},\vec{k}}=\frac{\rm sechr_{k}(\eta)}{1+\mu_{2}\tanh r_{k}(\eta)}\sum_{n=0}^{\infty}|1-\mu_{1}^{2}|^{\frac{n}{2}}\frac{(-\exp(2i\phi_{k}(\eta))\tanh r_{k}(\eta))^{n}}{(1+\mu_{2}\tanh r_{k}(\eta))^{n}}|n,n\rangle_{-\vec{k},\vec{k}},$ (15)
where $|n,n\rangle_{-\vec{k},\vec{k}}$ is the Krylov basis in this work, $\eta$ is the conformal time, $r_{k}$ and $\phi_{k}$ are two squeezing parameters, $\mu_{2}$ and $|1-\mu_{1}^{2}|$ as defined in Eq. (13). Here, we will introduce more information on wave function (15), in which this wave function will describe the quantum fluctuations in inflation. First, this wave function could represent the dissipative effects of one dynamical system denoted by $\mu_{2}$, and it will nicely recover the two-mode squeezed state in the weak dissipative approximation $(\mu_{2}\ll 1)$,
$|\mathcal{O}(\eta))=\frac{1}{\cosh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}e^{2in\phi_{k}}\tanh^{n}r_{k}|n;n\rangle_{\vec{k},-\vec{k}}+\mathcal{O}^{n}(\mu_{2}).$ (16)
Thus, we can see that the pure two-mode squeezed state is a wave function that corresponds to a closed system. In later illustrations, we will confirm this point. Secondly, the derivation of the OTMSS (15) only relies on the second kind of Meixner polynomial [48],
$P_{n+1}(x)=(x-\tilde{c}_{n})P_{n}(x)-b_{n}^{2}P_{n-1}(x),$ (17)
where we define $\tilde{c}_{n}=-ic_{n}$, and $x$ denotew the Hamiltonian. The sequence begins with the initial conditions $P_{0}(x)=1$ and $P_{1}(x)=x-c_{0}$. The polynomial $P_{n}(x)=\det(x-\mathcal{L}_{n})$ corresponds to an open system, and $\mathcal{L}_{n}$ represents the Liouvillian superoperator for the $n$-th quantum state. By introducing the natural orthonormal basis $\{e_{n}\}$. We could represent $|P_{n}(x))=\Bigl{(}\prod_{k=1}^{n}b_{k}\Bigr{)}|\mathcal{O}_{n}),\ \ \mbox{and}\ \ \ |x^{n})=\mathcal{L}^{n}|\mathcal{O}).$ Combining Eq. (17), there is an explicit relation as follows,
$b_{n+1}e_{n+1}+b_{n}e_{n-1}=(x-\tilde{c}_{n})e_{n},$ (18)
where $e_{n}=\mathcal{O}_{n}$ in our case and then it is equivalent to the generalized Lanczos algorithm (6). Through this simple illustration, one could explicitly see that the OTMSS only relies on the second kind of Meixner polynomial, describing the dissipative effects of open system (the early universe). Therefore, we could see that the OTMSS (15) is more generic compared with the pure two-mode squeezed state. Finally, we need to emphasize that the OTMSS is model-independent since the information of various models is encoded in $\mu_{2}$ and $|1-\mu_{1}^{2}|$.
V Bogoliubov transformation
Since the power spectrum of BD vacuum is well consistent with observations [6], we will use the Bogoliubov transformation to construct the relation between the BD vacuum and the vacuum of (15). First, the mode function $v_{z}(\eta,\vec{k})$ parametrizes the squeezed vacuum state in time-evolving spacetime. As for the reference BD vacuum mode, it is denoted by $v_{BD}(\eta,\vec{k})$. Then, we could introduce the non-trivial Bogoliubov transformation to relate $v_{z}(\eta,\vec{k})$ and $v_{BD}(\eta,\vec{k})$ as follows,
$v_{z}(\eta,\vec{k})=\alpha_{k}v_{BD}(\eta,\vec{k})+\beta_{k}v_{BD}^{*}(\eta,\vec{k}),$ (19)
where the coefficient of Bogoliubov transformation satisfying with the canonical Wronskian condition,
$|\alpha_{k}|^{2}-|\beta_{k}|^{2}=1.$ (20)
Meanwhile, the annihilation operator of OTMSS can also be defined by
$\hat{c}_{\vec{k}}=\alpha_{k}\hat{a}_{\vec{k}}+\beta_{k}^{*}\hat{a}_{-\vec{k}}^{\dagger}.$ (21)
Then, the we could use the Bogoliubov coefficient to construct the wave function via BD vacuum as follows,
$|0,0\rangle_{OST}=\frac{1}{\mathcal{N}}\exp\bigg{[}-\int\frac{d^{3}k}{(2\pi)^{3}}\frac{1}{2}\frac{\beta_{k}^{*}}{\alpha_{k}^{*}}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\bigg{]}|BD\rangle,$ (22)
where $\mathcal{N}$ represents the normalization factor, and $|0,0\rangle_{OST}$ denotes the vacuum of the OTMSS. The essence of Eq. (22) is that one can construct the wave function in the target spacetime using the Bogoliubov transformation and the BD vacuum. Therefore, the key aspect is determining the formulas for $\alpha_{k}$ and $\beta_{k}$ for our specific case. This is also crucial for understanding the relationship between the modes of the squeezed state and the modes of the BD vacuum, as indicated in Eq. (23). In fact, we can apply the squeezed operator $\mathcal{\hat{S}}_{-k,k}=\exp\left(\frac{r_{k}}{2}e^{-i\phi_{k}}\hat{a}_{k}\hat{a}_{-k}-\frac{r_{k}}{2}e^{i\phi_{k}}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\right)$ to the annihilation operator of the BD vacuum to derive the coefficients for the Bogoliubov transformation,
$\mathcal{\hat{S}}^{\dagger}\hat{a}\mathcal{\hat{S}}=\hat{c}_{\vec{k}}=\cosh r_{k}\hat{a}_{\vec{k}}-e^{i\phi_{k}}\sinh r_{k}\hat{a}_{-\vec{k}}^{\dagger},$ (23)
Compared with Eq. (21), we can obtain the formula for $\alpha_{k}$ and $\beta_{k}$
$\alpha_{k}=\cosh r_{k},~{}~{}~{}\beta_{k}=-\exp(-i\phi_{k})\sinh r_{k},$ (24)
where $\phi_{k}$ and $r_{k}$ are two parameters depicting the properties for the OTMSS (15) and the formula of $\alpha_{k}$ and $\beta_{k}$ is consistent with [49, 50]. Next, we will proceed with the evolution of $\phi_{k}$ and $r_{k}$ determined by the Hamiltonian (11).
As we discussed, the OTMSS (15) is more generic compared with a pure two-mode squeezed state, including the dissipative effects of an open system. In this work, we will particularly focus on the super Hubble scale, namely, $k\eta$ will be approaching zero. In light of (13), we could see that $\mu_{2}$ (dissipative coefficient) will also be very tiny in this super Hubble scale, meaning that the universe within the super Hubble scale is a weak dissipative system. Consequently, the pure two-mode squeezed state will be a good approximation, and our derivation for (24) is also consistent with each other.
VI Evolution of $\phi_{k}$ and $r_{k}$
To obtain the power spectrum of the OTMSS (15), one needs to rewrite the Bogoliubov coefficients (24) in terms of $k$ as in Refs. [51, 52, 53], where their formulas has made some approximation. To fully capture the information of (15), we will explicitly numerically give the evolution of $r_{k}$ and $\phi_{k}$ in terms of $k$, where the evolution is completely determined by the Schrödinger equation $\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle$. The resulting evolution equation of $r_{k}$ and $\phi_{k}$ can be denoted by via [26]
$r_{k}^{\prime}=\frac{-M_{P}|1-\mu_{1}^{2}|\sinh 2r_{k}\cos 2\phi_{k}-\sinh 2r_{k}\cdot\mu_{2}^{\prime}}{\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k}},$ (25)
$\phi_{k}^{\prime}=-M_{P}\mu_{2}+\frac{1}{2}M_{P}|1-\mu_{1}^{2}|^{\frac{1}{2}}\sin 2\phi_{k}\Big{[}|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\coth r_{k}+\mu_{2})\Bigr{]}.$ (26)
Compared to the original evolution equations for $\phi_{k}$ and $r_{k}$, we have neglected the contribution from the potential due to the slow-roll conditions. These calculations can be quite complicated; for more details, the reader can refer to the appendix of [26]. The evolution equations (26) for $r_{k}$ and $\phi_{k}$ describe the dynamics during the inflationary period. The power spectrum was observed at the horizon crossing at a fiducial momentum of $k_{*}=0.05~{}\text{Mpc}^{-1}$. To compare with observations, the data from the CMB covers the range from $10^{-4}~{}\text{Mpc}^{-1}$ to $1~{}\text{Mpc}^{-1}$ [6], making it beneficial to fully expand the power spectrum within this range. At the pivot point, we have $k_{*}=aH$, where the value is derived from today’s measurements, allowing us to easily obtain $\eta=-\frac{1}{k}$ within the CMB observational range [54, 55, 56]. Given that $\eta=-\frac{1}{aH}$, the evolution can be transformed accordingly,
$\begin{split}r_{k}^{\prime}=&\ -\frac{\tanh r_{k}(\mu_{2}^{\prime}+M_{P}|1-\mu_{1}^{2}|\cos 2\phi_{k})}{\tanh r_{k}+\mu_{2}},\\ \phi_{k}^{\prime}=&\ -M_{P}\mu_{2}+\frac{1}{2}\sin 2\phi_{k}\Bigl{(}\frac{M_{P}|1-\mu_{1}^{2}|\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\\ &\ +M_{P}\coth r_{k}+M_{P}\mu_{2}\Bigr{)}.\end{split}$ (27)
Being armed with these two equations, we could give their numeric in terms of comoving momentum $k$.
Figure 1: The numeric of squeezed parameter $r_{k}$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
Figure 2: The numeric of squeezed parameter $\phi_{k}$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
From Fig. 1 and Fig. 2, one can see that the evolution of $r_{k}$ and $\phi_{k}$ are almost the constants from $10^{-4}~{}Mpc^{-1}$ to $1~{}Mpc^{-1}$. Even, we can further evaluate the number density of OTMSS in terms of BD vacuum,
$\langle 0,0|\hat{c}^{\dagger}_{\vec{k}}\hat{c}_{-\vec{k}}|0,0\rangle=|\beta_{k}|^{2}=\sinh^{2}r_{k},$ (28)
where $|0,0\rangle$ is the BD vacuum and we have implemented Eq. (21). In light of this equation, we can give the numeric of mean number density under the BD vacuum.
Figure 3: The numeric of squeezed parameter $|\beta_{k}|^{2}$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
Fig. 3 clearly indicates that mean number density is almost zero since the particles are generated in preheating process, although the number density of particle is non-vanishing. Being armed with the evolution of $r_{k}$ and $\phi_{k}$ in terms of k, one can investigate the power spectrum.
VII Power spectrum of OTMSS
Since the power spectrum of OTMSS (15) is proportional to $|v_{z}|^{2}$. Then, we can obtain the ratio between the power spectrum of the OTMSS and the BD vacuum in light of (23),
$\gamma_{z}=\frac{\Delta_{\mathcal{R}z}^{2}}{\Delta_{\mathcal{R}}^{2}}=|\alpha-\beta|^{2}=\cosh 2r_{k}+\sinh 2r_{k}\cos\phi_{k},$ (29)
where $\Delta_{\mathcal{R}z}^{2}$ denotes the power spectrum for the OTMSS, $\Delta_{\mathcal{R}}^{2}$ represents the power spectrum of BD vacuum and we have used $\lim_{k\eta\to 0}\frac{v_{BD}^{*}}{v_{BD}}=-1$ (super Hubble scale approximation) via BD vacuum (5). Once having this relation, it is quite straightforwardly showing the numeric of (29) in terms of k.
Figure 4: The numeric of squeezed parameter $\gamma_{z}(k)$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
Fig. 4 indicates that the ratio (29) is almost one after the horizon crossing, inferring that the power spectrum of two-mode squeezed state will be consistent with the BD vacuum. For completeness, we will use the definition of power to confirm this conclusion. Following the notation of [34], the two-point correlation function can be defined as
$\begin{split}\langle|\hat{v}_{z}|^{2}\rangle=&\ \langle 0,0|\hat{v}_{z}^{\dagger}(\eta,0)\hat{v}_{z}(\eta,0)|0,0\rangle\\ =&\ \int d\ln{k}\frac{k^{3}}{2\pi^{2}}|v_{BD}(\eta,\vec{k})|^{2}\gamma_{z}.\end{split}$ (30)
where $\gamma_{z}$ is the ratio (29). Thereafter, we could define the power spectrum of the mode function $v_{z}$ as follows,
$\begin{split}\Delta_{v_{z}}^{2}(\eta,k)=&\ \frac{k^{3}}{2\pi^{2}}|v_{BD}(\eta,\vec{k})|^{2}\gamma_{z}\\ =&\ \Delta_{v_{BD}}^{2}(\eta,k)\gamma_{z},\end{split}$ (31)
To match the observations, we will use the curvature perturbation to calculate the power spectrum, where the curvature perturbation is $\mathcal{R}=v/z$ with $z=\sqrt{2\epsilon}a$ as shown in Sec. II. Thus, we could obtain the power spectrum of curvature perturbation as follows,
$\Delta_{\mathcal{R}z}^{2}(\eta,k)=\frac{1}{2\epsilon a^{2}}\frac{\Delta_{v_{z}}^{2}(\eta,k)}{M_{P}^{2}}=\frac{1}{2\epsilon a^{2}}\frac{\Delta_{v_{BD}}^{2}\gamma_{z}(\eta,k)}{M_{P}^{2}},$ (32)
As for the power spectrum of BD vacuum, it will predict the $\Delta_{\mathcal{R}}^{2}=A_{s}(\frac{k}{k_{*}})^{n_{s}-1}$ where the observational constraints to the amplitude power spectrum $A_{s}$ and spectral index from [6]
$A_{s}=(2.196\pm 0.060)\times 10^{-9},~{}~{}~{}n_{s}=0.9649\pm 0.0042.$ (33)
Based on these observational constraints, we could give the prediction for the power spectrum of OTMSS in light of (32).
Figure 5: The numeric of squeezed parameter $\Delta_{\mathcal{R}z}^{2}(k)$ in terms of $k$ in the horizon crossing. The dotted line is the reference point of $k_{*}=0.05\ \mbox{Mpc}^{-1}$.
Fig. 5 clearly shows the power spectrum of the OTMSS is almost the same with BD vacuum due to the ratio relation (29) as shown in Fig. 4, meanwhile our evaluation for the spectral index $n_{s}$ is also consistent with the BD vacuum.
Recalling the calculation of the power spectrum for the squeezed state as outlined in [51, 52, 53], we can parametrize the Bogoliubov coefficients, $\alpha_{k}$ and $\beta_{k}$, in terms of $r_{k}$ and $\phi_{k}$. These coefficients will be expressed as functions of the momentum $k$. However, these procedures overlook a crucial aspect: both $r_{k}$ and $\phi_{k}$ evolve over time as determined by the corresponding Schrödinger equation, given by $\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle$. Therefore, a more accurate approach would be to take the time evolution of $r_{k}$ and $\phi_{k}$ into consideration. Furthermore, we focus on the OTMSS, as described in (15), which represents the open system (the universe is regarded as an open system). Following this approach, our numerical results for the power spectrum are consistent with the BD vacuum. While our derivation cannot be distinguished from the BD vacuum, the wave function in (15) is more general, being model-independent and relying solely on the second Meixner polynomial and the Lanczos algorithm.
VIII Conclusion and Discussion
Under the single field inflation, we have constructed a framework for calculating the power spectrum of two mode squeezed state, where the key ingredient is the OTMSS (15) and the generalized Lanczos algorithm as shown in Sec. III.
In light of the Lanczos algorithm, the Hamiltonian can be separated into components corresponding to open and closed systems, as shown in Eq. (14). During inflation, an amplification process drives quantum phenomena toward classicalization, making the two-mode squeezed state a natural framework for studying curvature perturbations. The squeezed state is constructed via the action of the squeezed operator $\mathcal{\hat{S}}_{-k,k}=\exp\left(u\hat{a}_{k}\hat{a}_{-k}-\bar{u}\hat{a}_{k}^{\dagger}\hat{a}_{-k}^{\dagger}\right)$ on the BD vacuum, where $u=\frac{r_{k}}{2}e^{-i\phi_{k}}$ in our case. Crucially, (14) reveals that $\mathcal{\hat{S}}_{-k,k}$ exclusively encodes the closed system dynamics. In contrast, the OTMSS in (15) incorporates environmental interactions through the dissipative coefficient $\mu_{2}$. The standard squeezed state emerges as the leading-order term in the weak dissipation limit ($\mu_{2}\to 0$). From the Schrödinger equation $\hat{H}|\mathcal{O}(\eta)\rangle=i\partial_{\eta}|\mathcal{O}(\eta)\rangle$, we derive the evolution equations for $r_{k}(\eta)$ and $\phi_{k}(\eta)$, whose numerical solutions are plotted in Figs. 2 and 1. The ratio between the OTMSS and BD vacuum, defined via Bogoliubov transformations in Eq. (29). As shown in Fig. 4, $\gamma_{k}$ remains unity throughout inflation, consistent with suppressed particle production (mean particle density $\langle n_{k}\rangle\ll 1$) due to preheating dynamics. Consequently, the power spectrum $\Delta_{\mathcal{R}}^{2}(k)$ computed from the OTMSS in (15) matches the BD vacuum prediction, as demonstrated in Fig. 5.
Key insights:
  • Dissipation ($\mu_{2}$) modifies higher-order correlations in the strong dissipative regime but leaves the leading-order power spectrum intact.
  • Observational indistinguishability from BD vacuum arises from $\gamma_{k}\approx 1$, even with $\mu_{2}\neq 0$.
  • Lanczos-based separation of $\hat{H}$ clarifies the open system’s part and closed system’s part.
The derivation of the BD vacuum in (5) relies on solving the mode equation. In contrast, the OTMSS is constructed through the Hamiltonian formalism and Lanczos algorithm without solving equations of motion, where model-dependent information is encoded in the parameters $\mu_{2}$ and $|1-\mu_{1}^{2}|$. Remarkably, the power spectrum prediction from the OTMSS in (15) matches the BD vacuum result, suggesting that squeezed state construction is a more fundamental approach than solving EOM.
The key ingredient in this construction is the generalized displacement operator [57], defined as
$\mathcal{D}(\xi)=\exp\left(\xi\hat{H}-\bar{\xi}\hat{H}^{\dagger}\right),$ (34)
where $\hat{H}$ exhibits specific group structures (e.g., the Weyl algebra in our case) and the squeezed operator belongs this kind of operator. When $\hat{H}$ belongs to a Lie group, the operator $\mathcal{D}(\xi)$ provides a rigorous framework for generating quantum states, surpassing traditional EOM-based methods in universality.
Extensions and implications:
  • Current results apply to single-field inflation. Generalizations to multifield scenarios [58, 59, 60, 61, 62] and $f(R)$ gravity [63, 64] are natural next steps.
  • The $\mu_{2}$-dependence in OTMSS may encode decoherence effects observable in higher-order correlations.
  • Group-theoretic construction offers a unified approach for quantum fluctuations in curved spacetime.
Acknowledgements.
LHL and KHZ are funded by National Natural Science Foundation of China (NSFC) with grant NO. 12165009, Hunan Natural Science Foundation with grant NO. 2023JJ30487 and NO. 2022JJ40340. HQZ is funded by NSFC with grant NO. 12175008.